SAVING (S) INCOME (Y) -1000 0 -500 2000 0 4000 500 6000 1000 8000 3. Answer questions for the table above, where the equation has the form S = a + bY.

a. What is the value of the vertical intercept?

b. What is the equation of the line?

c. This function exhibits what type of relationship?
Answer #1

Answer

(a)

Here Vertical axis is S and Horizontal axis is Y.

Vertical intercept is the value of S at which Y = 0. We can see from above that when Y = 0, S = -1000

Hence, Vertical Intercept = -1000

(b)

Equation of straight line is of the form :

(y - yo) = (Change in y / Change in x)(x - xo)

Take two points (Y,S) = (0,-1000) and (4000, 0)

Change in y = 0 - (-1000) = 1000 and Change in x = 4000 - 0 = 4000

Hence (Change in y / Change in x) = 1000/4000 = 0.25

Hence S - (-1000) = 0.25(Y - 0)

=> S = -1000 + 0.25Y ---------------------Saving function

(c)

We can see from above that as Y increases S will also increase. This means that there is direct relation between Y and S and the above function is called Saving function, where Autonomous saving = -1000 and Marginal Propensity to consume (MPC) = 0.25